# Private Fine-tuning of Large Language Models with Zeroth-order Optimization

### v2

@@ Appendix B Implementation Details
We follow Malladi et al. [2023] and provide the memory-efficient version of DP-ZO in Algorithm 2. Algorithm 2 enjoys the benefit that it does not incur additional GPU memory cost compared to inference.
Algorithm 2 Differentially Private-ZO (GPU memory efficient version. Adapted from Malladi et al. [2023])
@@ Appendix C Design Choices
Algorithm 1 estimates the gradients once. As outlined above, SPSA can be extended to n-SPSA gradient estimator and n-SPSA can improve the performance in the non-private setting [Malladi et al., 2023]. Here we discuss our design choice of why we choose $n=1$ in Algorithm 1.
  • Estimate the average. Previous work [Malladi et al., 2023] shows that averaged estimation helps the non-private setting. In a private setting, we have to privatize the gradient estimation. Here we discuss our initial design of the privatized n-SPSA gradient estimation. For the sampled batch, assuming we are adding the Gaussian noise $\mathcal{N}(0,\mathrm{C}^{2}\sigma^{2})$ for 1-SPSA. Then for n-SPSA, to ensure we have the same privacy cost as 1-SPSA, we need to add $\mathcal{N}(0,n\cdot\mathrm{C}^{2}\sigma^{2})$ to each gradient estimation and finally average the $n$ gradients. Our privacy analysis follows the $n$-fold composition of Gaussian mechanism (Corollary 3.3 in Gaussian differential privacy [Dong et al., 2019]). Our initial experiment result shows that our current analysis for n-SPSA noise addition does not make n-SPSA improve in the private setting compared to 1-SPSA. We leave the improvement in tighter analysis for private n-SPSA as future work and use 1-SPSA to conduct experiments.
@@
As discussed in Section 3Algorithm 1 can be incorporated in different noise mechanisms. We focus on the Gaussian noise mechanism and the Laplace mechanism in this work. The Gaussian noise mechanism has been widely studied in previous literature both for privacy analysis and empirical performance in DP-SGD [Abadi et al., 2016, Mironov, 2017, Dong et al., 2019]. The Laplace mechanism, though less studied for privacy-preserving machine learning, can provide pure DP while the Gaussian mechanism can only provide approximate DP. We have provided the privacy analysis in Section A.
Appendix D Hyperparameter Search
Our experiments are based on the open-source code [^2] https://github.com/princeton-nlp/MeZO. of Malladi et al. [2023]. We provide the prompts we use in Table Table 13. In this section, we first provide several initial results for hyperparameter search on clipping threshold and finally present the hyperparameter tables. We also provide an initial study to systematically evaluate the interplay between batch size and training iterations for DP-ZO.
@@ Appendix D Hyperparameter Search
Li et al. [2022b], De et al. [2022] recommend small clipping $C$ threshold for DP-SGD training. For example, Li et al. [2022b] use $C=0.1$ for training language models. We therefore study the effect of different clipping threshold and present the results in Table 14. We find that while $C=1$ performs significantly worse, setting $C$ as $0.1$, $0.05$, $0.01$ are within the 2% performance gap. We therefore choose $C=0.05$.
@@
Hyperparameter for DP-ZO (Gaussian) in Main Results.
We present the hyperparameter for DP-ZO (Gaussian) in Table 15 and Table 16.
Table 15: Hyperparameter search for DP-ZO in main results Table 1.
@@
Hyperparameter for DP-SGD.
We present the hyperparameter search for DP-SGD in Table 17.
Table 17: Hyperparameter search for DP-SGD in Table 2.
@@
Hyperparamter for DP-ZO (Laplace).
The hyperparameter search for DP-ZO (Laplace) is similar to DP-ZO (Gaussian).
Ablation on Batch Size and Steps.
In Table 19 and Table 19, we did an initial study to systematically evaluate the interplay between batch size and training iterations by varying batch size in [16,32,64,128] and steps in [10000, 2000, 40000, 80000]. Similar to main results, we run 5 independent runs for each setting and compute the average of 5 runs. This ablation is by OPT-13B on SQuAD dataset with LoRA fine-tuning. Table 19 and Table 19 show that increasing steps $T$ improves the performance more than increasing the batch size. We also did ablation study on $T$ in [200, 400, 800, 1600] for DP-SGD (and did not observe significant improvements in DP-SGD) to ensure the fair comparison of DP-SGD and DP-ZO. Taking the computation limitation into consideration, we set $T=75000$ and BSZ=16 for main results in Table 1. We leave more investigation on the batch size and steps for DP-ZO, such as variance reduction method, as future work.
Table 18: $T=10000$, Varying batch size.
@@
Appendix E Ablation on Model Size
@@
Table 20: Ablation of DP-ZO across different model sizes. $(1,10^{-5})$-DP.

### v3

@@ Appendix A Privacy Analysis
Approximate $\varepsilon$-DP by Laplace mechanism. We can also get tighter composition of $\varepsilon$ with relaxation to $\delta>0$. The most straight forward way is to instantiate the privacy loss random variable of random response with $(\log(1+p\cdot(e^{1/\sigma}-1)),0)$ because the dominating pair for random response is a dominating pair for the pure DP mechanism. Then, we can use the numerical composition of Pure DP PRV accountant by Gopi et al. (2021). Note that this method is agnostic to the DP mechanisms used for pure $\varepsilon$-DP. We now provide a more fine-grained privacy analysis for the Laplace mechanism. Specifically, we could compute the privacy cost of composition for the Laplace mechanism by Monte Carlo based DP accountant (Wang et al., 2023). Note that since we are dealing with scalar values, our mechanism in each iteration will be a one dimensional Laplace mechanism. Let $b$ be the scale of Laplace noise, $p$ the sub-sampling rate, and assume the sensitivity is $1$, and assume we are doing composition for $T$ iterations, each iteration with sampling rate $p$. By Zhu et al. (2022) we know that the pair of distribution $(P,Q)$ dominating pair for a single dimensional Laplace mechanism, where $P$ and $Q$ are distributed according to the following probability density functions,
Therefore, ($P,(1-p)\cdot P+p\cdot Q$) is the dominating pair for the sub-sampled Laplace. We plug this into the standard Monte-Carlo accountant of Wang et al. (2023) (without importance sampling, see Algorithm 2 and Theorem 10 in Wang et al. (2023)) while using $10^{10}$ samples to calculate the $\delta$ at a given value of $\epsilon$. Also, using the analytical accountant explained above, we always make sure that $\mathbb{E}[\hat{\delta}_{MC}^{2}]$ is bounded by $10^{-8}$ (we use the fact that $\mathbb{E}[\hat{\delta}_{MC}^{2}]$ is bounded by $\mathbb{E}[Y^{2}]$ and the fact the PRV $Y$ is always bounded for Laplace mechanism). This ensures that the error of our estimation of $\delta$ is at most $10^{-8}$ with probability at least $1-10^{-5}$. Putting all together, for all reported values of $\varepsilon$, our $\delta$ is bounded by $10^{-5}$, with probability at least $0.99999$. This privacy analysis is tighter with $\varepsilon$ is high compared to the former analysis which uses the pure DP PRV accountant. This is consistent with the intuition. As we increase the distance between the Laplace dominating pairs, the probability of sampling points from the area between the centers increases. And that is where the Laplace Mechanism is different from the Randomized Response. We present the accounting results for the Laplace method to achieve $(\varepsilon,\delta)$-DP by these two accounting methods in Table 13. Table 13 shows that the Monte Carlo based DP accountant can give tighter analysis for the Laplace mechanism for $(\varepsilon,\delta)$-DP than the pure $\varepsilon$-DP PRV method.
Remark A.9.
We want to note a minor technicality in our privacy analysis with poisson subsampling. Zhu et al. (2022) treat the add neighboring relation and remove neighboring relation separately, because this is crucial in retaining a tight dominating pair with a closed-form expression. As shown in Zhu et al. (2022) and Lebeda et al. (2025), neither one of the add/remove neighboring relation can dominates the other one. To get the results for the standard add/remove neighboring relation for a $T$-fold composition of subsampled mechanism, we need to do the pointwise maximum of the add-only and remove-only relation. The Microsoft (2021) library only implements one of the add/remove and our randomized accounting method for subsampled Laplace mechanism only considers the add-only relation. We note that Google’s-DP-Library (2020) implements DP accounting by taking maximum of the two neighboring relations under $T$ folds and provides the accounting code for both sub-sampled Gaussian and sub-sampled Laplace mechanisms. We verify that our privacy parameters in Tables 13, 16, 17 and 18 do not exceed the desired $\varepsilon$ using the method of Doroshenko et al. (2022) in Google’s-DP-Library (2020) with the pessimistic estimate (i.e., the upper bound).
@@ Appendix B Implementation Details
We follow Malladi et al. (2023) and provide the memory-efficient version of DP-ZO in Algorithm 2. Algorithm 2 enjoys the benefit that it does not incur additional GPU memory cost compared to inference.
Algorithm 2 Differentially Private-ZO (GPU memory efficient version. Adapted from Malladi et al. (2023))
@@ Appendix C Design Choices
Algorithm 1 estimates the gradients once. As outlined above, SPSA can be extended to n-SPSA gradient estimator and n-SPSA can improve the performance in the non-private setting (Malladi et al., 2023). Here we discuss our design choice of why we choose $n=1$ in Algorithm 1.
  • Estimate the average. Previous work (Malladi et al., 2023) shows that averaged estimation helps the non-private setting. In a private setting, we have to privatize the gradient estimation. Here we discuss our initial design of the privatized n-SPSA gradient estimation. For the sampled batch, assuming we are adding the Gaussian noise $\mathcal{N}(0,{C}^{2}\sigma^{2})$ for 1-SPSA. Then for n-SPSA, to ensure we have the same privacy cost as 1-SPSA, we need to add $\mathcal{N}(0,n\cdot{C}^{2}\sigma^{2})$ to each gradient estimation and finally average the $n$ gradients. Our privacy analysis follows the $n$-fold composition of Gaussian mechanism (Corollary 3.3 in Gaussian differential privacy (Dong et al., 2019)). Our initial experiment result shows that our current analysis for n-SPSA noise addition does not make n-SPSA improve in the private setting compared to 1-SPSA. We leave the improvement in tighter analysis for private n-SPSA as future work and use 1-SPSA to conduct experiments.
@@
As discussed in Section 3Algorithm 1 can be incorporated in different noise mechanisms. We focus on the Gaussian noise mechanism and the Laplace mechanism in this work. The Gaussian noise mechanism has been widely studied in previous literature both for privacy analysis and empirical performance in DP-SGD (Abadi et al., 2016; Mironov, 2017; Dong et al., 2019). The Laplace mechanism, though less studied for privacy-preserving machine learning, can provide pure DP while the Gaussian mechanism can only provide approximate DP. We have provided the privacy analysis in Section A.
Appendix D Experimental Details
D.1 Datasets and Metrics
Datasets. Following Malladi et al. (2023), we use SST2 (Socher et al., 2013) for text classification and SQuAD (Rajpurkar et al., 2016) and DROP (Dua et al., 2019) for text generation. SST2 is a binary classification for sentiment classification based on text from movie reviews. SQuAD and DROP are question answering tasks, that given the context and question, the language model should output needed answers. Such answers are generated tokens one by one and therefore considered as generation task. In Section 4.3, we follow prior works (Yu et al., 2022; Li et al., 2022b) and run experiments on QNLI (Wang et al., 2019). QNLI is a binary classification task to determine whether the two sentences in a pair are entailment or not.
[…]
D.2 Hyper-parameter Search
Our experiments are based on the open-source code [^4] https://github.com/princeton-nlp/MeZO. of Malladi et al. (2023). We provide the prompts we use in Table 14. In this section, we first provide several initial results for hyperparameter search on clipping threshold and finally present the hyperparameter tables. We also provide an initial study to systematically evaluate the interplay between batch size and training iterations for DP-ZO.
@@ D.2 Hyper-parameter Search
Li et al. (2022b); De et al. (2022) recommend small clipping $C$ threshold for DP-SGD training. For example, Li et al. (2022b) use $C=0.1$ for training language models. We therefore study the effect of different clipping threshold and present the results in Table 15. We find that while $C=1$ performs significantly worse, setting $C$ as $0.1$, $0.05$, $0.01$ are within the 2% performance gap. We therefore choose $C=0.05$.
@@
Hyper-parameter for DP-ZO (Gaussian) in Main Results.
We present the hyper-parameter for DP-ZO (Gaussian) in Table 16 and Table 17.
Table 16: Hyper-parameter search for DP-ZO in main results Table 1.
@@
$\phi$ 0.001
Hyper-parameter for DP-SGD.
We present the hyper-parameter search for DP-SGD in Table 18.
Table 18: Hyper-parameter search for DP-SGD in Table 2.
@@
Hyper-parameter for DP-ZO (Laplace).
The hyper-parameter search for DP-ZO (Laplace) is similar to DP-ZO (Gaussian).
Effects of Batch Size and Steps.
In Table 20 and Table 20, we did an initial study to systematically evaluate the interplay between batch size and training iterations by varying batch size in [16,32,64,128] and steps in [10000, 2000, 40000, 80000]. Similar to main results, we run 5 independent runs for each setting and compute the average of 5 runs. This set of experiments is done on OPT-13B on SQuAD dataset with LoRA fine-tuning. Table 20 and Table 20 show that increasing steps $T$ improves the performance more than increasing the batch size. We also run experiments of $T$ in [200, 400, 800, 1600] for DP-SGD (and did not observe significant improvements in DP-SGD) to ensure the fair comparison of DP-SGD and DP-ZO. Taking the computation limitation into consideration, we set $T=75000$ and batch size BSZ=16 for main results in Table 1. We leave more investigation on the batch size and steps for DP-ZO, such as variance reduction method, as future work.
Table 19: $T=10000$, Varying batch size BSZ.
@@
Appendix E Effect of Model Size
@@
Table 21: The experiment results of DP-ZO across different model sizes. $(1,10^{-5})$-DP.
@@
Appendix F RELATED WORK
Privacy preserving large language models. Recent studies have leveraged DP-SGD to fine-tune large language models. Li et al. (2022b) provide methods for fine-tuning large language models with DP-SGD by ghost clipping to mitigate the memory burden of per-sample gradient clipping. Yu et al. (2022) report compelling results by only updating a sparse subset of the LLMs with parameter efficient fine-tuning (PEFT) methods such as LoRA (Hu et al., 2022). He et al. (2023) leverage group-wise clipping with adaptive clipping threshold and privately fine-tune the 175 billion-parameter GPT-3. Duan et al. (2023); Li et al. (2022b) also consider private prompt tuning by adding noise to the soft prompt (Li & Liang, 2021; Lester et al., 2021). Du et al. (2023) add non-i.i.d. noise from a matrix Gaussian distribution to directly perturb embedding in the forward pass of language models. With the emergence in-context learning of large language models (Brown et al., 2020), recent works (Duan et al., 2023; Wu et al., 2024; Tang et al., 2024) study private in-context learning of large language models without fine-tuning.
Private zeroth-order optimization. Most recently, a concurrent work (Zhang et al., 2024a) also considers the same DP-SPSA algorithm for zeroth-order optimization. We have discussed our work and DPZero (Zhang et al., 2024a) in Section 5.